APPENDIX: DATA SOURCES

DELIVERIES

Aggregate deliveries on CBOT futures contracts each month were assembled from the CBOT Statistical Annuals for 1964 through 1985. From 1986 on, the data were published in the CBOT Statistical Supplements. Public Affairs at the KCBOT provided delivery data for the KCBOT wheat contracts between 1972–90. Deliveries on Comex contracts were taken from their annual Statistical Yearbooks, 1977–89.

The records of daily deliveries identified by location on CBOT contracts were provided by the Economic Analysis and Planning Department at the CBOT for the period May 1976 through September 1989. These data were used in the calculation of the timing of deliveries and to identify the location of deliveries when multiple locations were possible. The daily data did not always aggregate to the totals reported in the Statistical Annuals, although in most cases the differences were comparatively small. Obvious errors in either series were discarded. Daily delivery records for Kansas City wheat were provided by the KCBOT. Comex Statistical Yearbooks contain these data for copper.

FUTURES MARKET POSITIONS

Volume and open interest data were taken variously from CBOT Statistical Annuals, the Wall Street Journal (noting that WSJ reports are always “as of the previous day”), and daily computerized records provided by the Economic Analysis and Planning Department at the CBOT. Comex data were available in their Yearbooks. Volume and open interest data for the KCBOT wheat market were taken from the Kansas City Grain Market Review, a daily publication.

Data on the concentration of futures positions during the delivery month were provided by the CFTC. These data covered CBOT wheat, corn, and soybean contract expirations from 1982 through September 1989. The data are daily reports of the positions in the expiring contract of the four largest traders on both the long and short sides of the market for a period 5

Food Research Institute Studies, Vol. XXII, No. 2, 1991.

days before the first delivery day through the last trading day in the expiring contract month. The only exception is for January soybeans where the records for the five days prior to the first delivery day could not be included due to problems in the data records at the CFTC when crossing calendar years. The CFTC also provided concentration data for Kansas City wheat and Comex copper for the 1982–89 period. Inspection of their records from 1982, 1983, and 1984 revealed significant reporting problems, and the data appeared not to be very reliable. Thus, only the 1985–89 data were used in the comparative analyses.

PRICES

Cash prices examined in the analyses of the wheat, corn, and soybean markets are primarily USDA-reported price series for Chicago and Toledo. They were compiled from the USDA, Agricultural Marketing Service, Grain and Feed Market News (formerly Grain Market News), Weekly Summary and Statistics, 1964–89. Use was also made of the CBOT Statistical Annuals for reports of the same price series; however, the Annuals contained numerous errors, especially in more recent years. Neither the USDA nor the CBOT reported daily cash market prices before 1970. The Wall Street Journal and the Minneapolis Daily Market Record have been referred to for Chicago prices on the specific days required by the analyses of this period.

Gulf prices for hard and soft wheat, corn, and soybeans were also taken from the USDA publication. The series were “Export Offers Prompt or 30 Day Shipment, Delivered Port” in the Gulf and, when distinguished, the barge quote was used. If ranges for these prices were reported, these were recorded. Preliminary analyses indicated the mean of the range was as good a measure of value in most of the analyses. A separate series of cash corn prices, daily 1973–90, in Chicago were provided by CPC International, Inc. They are the best, posted Chicago cash price bids for #2 yellow corn, daily 1973–90, and are published in the New York Times. The Andersons provided weekly (Wednesdays) basis bid prices in Toledo, 1977–89, for wheat, corn, and soybeans and weekly FOB Toledo values for soybeans from 1985– 90. Cash prices for wheat in Kansas City were taken from the Kansas City Grain Market Review.

Weekly barge freight rates on the Upper Illinois River were provided by Consolidated Grain and Barge Company, and when combined with Gulf prices, the rates permitted construction of an export value series for Chicago. The University of Illinois provided its weekly record of corn and soybean prices at some 13 separate elevator locations in the state. The records date to 1967 for the most part, and those from eight locations were judged to be complete enough to permit detailed analysis of their relations to futures prices.

STOCKS

Visible supplies are total stocks at selected terminal and elevator locations, reported weekly by the Agricultural Marketing Service of the USDA in “Stocks of Grain at Selected Terminal and Elevator Sites.” Until 1986, these data were also reported in the CBOT Statistical Annuals. The latter was used as the principal source, although care was taken to select the observations so that they were consistently for the Friday closest to the beginning of each month. From 1986–89, data were taken from the USDA publication directly. Quarterly national stocks are reported regularly in the USDA’s Wheat Situation and Feed Grain Situation; the USDA provided a convenient summary of the series needed.

Chicago free stocks of grain were constructed from various tables in the CBOT Statistical Annuals or Supplements. For 1964–69, the Annuals report “Free Stocks” directly. For 1970–89, the Annuals report “Total Stocks in Chicago” and “CCC Stocks in Chicago.” These tables, however, have consistently indicated that no CCC stocks were in Chicago (or Toledo) since the 1970s. In spite of their titles and associated notes, data in the tables in fact report free stocks in Chicago and Toledo, not totals. There were substantial quantities of CCC stocks in Toledo and Chicago during the 1980s, confirmed by the CFTC weekly “Stocks of Grain” reports. When total stocks in Chicago or Toledo were needed, resort was made to the CFTC and USDA stocks publications. Stocks of wheat in Kansas City were provided by the KCBOT or were found in the Grain Market Review; warehouse stocks of copper are in the Comex Statistical Yearbooks.